OpenStudy (anonymous):

I don't even know where to begin at, let alone understand how to do this.... Test the following equation for symmetry. Show all work. x=y^2-6

OpenStudy (amistre64):

f(-y) = f(y), its even; f(-y) = -f(y) its odd; if neither of those then its neither

OpenStudy (cwrw238):

first replace y with -y: x = (-y)^2 - 6 this is the same as x = y^2 - 6 so the graph is symmetrical about the x - axis

OpenStudy (cwrw238):

if we try replacing x with -x the above is not found so not symmetrical about y-axis

OpenStudy (dumbcow):

easy pick a y-value, say y = 2, then use the neg, y=-2 now see if they give you the same x value x = 2^2 -6 = -2 x = (-2)^2 -6 = -2